Work of 500 J is done by an ideal gas, when 500 cal of heat is added to the gas. What is the increase in the internal energy of the gas in J?

1
Expert's answer
2021-11-23T10:31:29-0500

"500\ cal=2092\ J"


"Q=\Delta U+W\to \Delta U=Q-W=2092-500=1592\ (J)" . Answer